Two point charges are located on the y-axis as follows: charge q1 = - 1.50 nC at y = - 0.600 m, and charge q2 = +3.20 nC at the origin (y = 0). What is the total force (magnitude and direction) exerted by these two charges on a third charge q3 = +5.00 nC located at y = -0.400 m?
1
Expert's answer
2020-04-13T10:06:09-0400

Given:- Two point charges on the y-axis

 q1 = - 1.50 nC  at ( y1 = - 0.600 m)

q2 = +3.20 nC at the origin (y2 = 0m)

q3 = +5.00 nC located at (y3= -0.400 m)

Find the total force (F) and (magnitude and direction)

Now,

"\\From\,Coulomb's\,law\\[5pt]
F=K\frac{|q_{1}q_{2}|}{r^{2}}\\[5pt]
The\,force\,exerted\,by\,the\,first\,charge\,on\,the\,third\,one\\[5pt]
r_{1}=y_{3}-y_{1}=-0.4m-(-0.6m)=0.2m\\[5pt]
\\The\,force\,is\, attractive\, in\, the\,negative \,y\,axis\\ 
F_{1}=K\frac{|q_{1}q_{2}|}{r^{2}}\\[5pt]
\\F_{1}=\frac{8.987\times 10^{9}N.m^{2}/C^{2}|(-1.5\times 10^{-9}C)(5\times 10^{-9}C)|}{(0.2m)^{2}}\\[5pt]
F_{1}=1.685\times 10^{-6}N\\[5pt]
Now,
\\The\,force\,exerted\,by\,the\,second\,charge\,on\,the\,third\,one,
\\r_{2}=y_{3}-y_{2}=-0.4m-0m=-0.4m\\[5pt]
\\The\,force\,is\, repulsive\, in\, the\,negative \,y\,axis\\ \\F_{2}=K\frac{|q_{2}q_{3}|}{r^{2}}\\[5pt]
F_{2}=\frac{8.987\times 10^{9}N.m^{2}/C^{2}|(3.2\times 10^{-9}C)(5\times 10^{-9}C)|}{(-0.4m)^{2}}\\[5pt]
F_{1}=8.99\times 10^{-7}N\\[5pt]
therefore,\\
Total\,force\,(F)\\[5pt]
\\F=F_{1}+F_{2}\\[5pt]
F=1.685\times 10^{-6}N+8.99\times 10^{-7}N\\[5pt]
F=2.59\times 10^{-6}N\\[5pt]
Direction:-(The \,negative\,y\,direction)"